Given: A= 004322111− B= 004322111− Find: a) –A^{-1}+ 3B^T b) B^{-1}+ ( A^T+A^{-1})


1
Expert's answer
2020-05-07T15:47:59-0400

"A= B =
 \left[ {\begin{array}{cc}
 0 & 0 & 4 \\
 3 & 2 & 2 \\
1 & 1 & 1
 \end{array} } \right]" .

By expanding along 1st row, "|A|=|B| = 4(3-2) = 4".

Now, "adj(A)=adj(B) = 
 \left[ {\begin{array}{cc}
 0 & -1 & 1 \\
 4 & -4 & 0 \\
-8 & 12 & 0 \\
 \end{array} } \right]^T = 
\left[ {\begin{array}{cc}
 0 & 4& -8 \\
 -1& -4 & 12 \\
1 & 0& 0 \\
 \end{array} } \right]" .

So,

"A^{-1} =B^{-1} = \frac{1}{|A|} adj(A) = \left[ {\begin{array}{cc}
 0 & 1 & -2\\
 -1/4 & -1& 3\\
1/4 & 0& 0 \\
 \end{array} } \right]" .

a)

"-A^{-1} + 3 B^T = - \left[ {\begin{array}{cc}
 0 & 1 & -2\\
 -1/4 & -1& 3\\
1/4 & 0& 0 \\
 \end{array} } \right]+ 3 \left[ {\begin{array}{cc}
 0 & 0 & 4 \\
 3 & 2 & 2 \\
1 & 1 & 1
 \end{array} } \right]^T 
\\
= \left[ {\begin{array}{cc}
 0 & -1 & 2\\
 1/4 & 1& -3\\
-1/4 & 0& 0 \\
 \end{array} } \right] + \left[ {\begin{array}{cc}
 0 & 9& 3\\
 0& 6& 3\\
12 & 6 & 3
 \end{array} } \right] = \left[ {\begin{array}{cc}
 0 & 8& 5\\
 1/4& 7& 0\\
47/4& 6 & 3
 \end{array} } \right]"

b)

"A^T + A^{-1} = \left[ {\begin{array}{cc}
 0 & 0& 4\\
 3& 2& 2\\
1 & 1 & 1
 \end{array} } \right] + \left[ {\begin{array}{cc}
 0 & 1 & -2\\
 -1/4 & -1& 3\\
1/4 & 0& 0 \\
 \end{array} } \right] = \left[ {\begin{array}{cc}
 0 & 1 & 2\\
 11/4 & 1& 5\\
5/4 & 1& 1\\
 \end{array} } \right]"